OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10

OpenStudy (welshfella):

Use the elimination method : - Multiply the first equation by 4 and the second equation by -3 The terms in y will then become 12y and -12y so adding the 2 equations eliminates y The result will be an equation in x

OpenStudy (welshfella):

- which you solve for x.

OpenStudy (anonymous):

can you do this with every single substitution problem?

OpenStudy (welshfella):

This is the elimination method Both the elimination and substitution method are used to solve systems of equations - the elimination method is usually the easiest one.

OpenStudy (anonymous):

after multiplying should i get 17x=43?

OpenStudy (welshfella):

8x + 12y = 4 9x - 12y = 30 adding:- 17x = 35

OpenStudy (welshfella):

I made an error in my first post - the second equation is multiplied by 3 not -3

OpenStudy (anonymous):

34?

OpenStudy (welshfella):

sorry 34

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

i have alot more questions would you mind answering

OpenStudy (welshfella):

yw - i dont know ehat wrong with me this morning - i keep on doing typos!!

OpenStudy (welshfella):

I have about 10 minutes so i can help with some

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18

OpenStudy (anonymous):

for this would you do the same and multiply line one by four and line two by 6

OpenStudy (anonymous):

is the correct asnwer 1

OpenStudy (welshfella):

- that would eliminate y and you need to find y

OpenStudy (anonymous):

i changed my steps i multiplied by negative 7 and positive 13 and i got possitive one is that correct

OpenStudy (welshfella):

- but you could do it that way - find x first - then y using substitution

OpenStudy (anonymous):

ohhhh that makes sense but is the answer positive 1

OpenStudy (welshfella):

lets see 91x - 42y = 140 91x + 52y = 234 add 8y = 374

OpenStudy (welshfella):

not sure if thats right - dont trust myself this morning

OpenStudy (anonymous):

the options were negative or positive one

OpenStudy (welshfella):

oh i;ll recheck

OpenStudy (welshfella):

ah - can you see my deliberate mistake ( lol 1)

OpenStudy (welshfella):

since the 2 terms in x are positive you need to multiply first by 13 and the second by -7 not 7

OpenStudy (welshfella):

91x - 42y = 140 -91x -52y = -234 add - 94y = -94 y = 1

OpenStudy (welshfella):

hope i've helped gotta go